OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4).

OpenStudy (cyber405):

If they are parallel then the slope of two line is the same... The equation of lines can be written like this... y=mx+b where m is slope of line and b is intercept of line. For example you can write 4x-3y=-12 like this...y=4/3x+4 The slope of this line is as same as the other...then you can write this y=4/3x+b then you can put the (-3,4) in the equation after solving it you have intercept which is b.

OpenStudy (cyber405):

Is that useful...?